# Force Prompting: Video Generation Models Can Learn and Generalize Physics-based Control Signals

### v1

@@ 6 Mass Understanding
Force-Mass Relationship Qualitative Study: We evaluate mass understanding across four benchmark tasks featuring geometrically identical objects with different implied masses. Our test scenarios includes ornaments (wooden versus cast iron), laundry baskets (empty versus filled with clothes), book stacks (one, two, or three books), and cube stacks (single versus double cube). To ensure experimental control, we utilize the GPT-Image-1 API to generate initial frames with variations where only the implied mass differs between conditions. Figure 6 presents some of these results, with demonstrating that lighter objects consistently travel farther when subjected to identical forces. This pattern remains robust across four random seeds. This behavior suggests an emergent understanding of mass-dependent physics in our force-prompted model. Other results are in the project webpage.
@@ Acknowledgments
We would like to thank Bill Freeman, Miki Rubinstein, Junyi Zhang, Junhwa Hur, Noah Fischer, Saining Xie, Calvin Luo, Shijie Wang, Koven Yu, Tian Yun, and Zilai Zeng for useful discussions. This project was partially supported by Samsung. Our research was conducted using computational resources at the Center for Computation and Visualization at Brown University.
@@ A.1 Training hyperparameters
We train for 5000 steps with an effective batch size of 8 (using gradient accumulation over two steps) and saved model checkpoints every five hundred steps. We used bf16 mixed precision with tf32 support and initialized from the THUDM/CogVideoX-5b-I2V pretrained weights. ControlNet (Zhang et al., 2023) was initialized from the first six transformer layers, for all three input channels, using a downscaling factor of eight, and a unit weight coefficient. We use the AdamW optimizer (Loshchilov and Hutter, 2019) (initial learning rate $1\times 10^{-5}$, $\beta_{1}=0.9$, $\beta_{2}=0.95$, maximum gradient norm 1.0) under a cosine-with-restarts learning-rate schedule (one cycle, 250 warm-up steps). We set our random seed to 42. Training videos contained up to 49 frames at 8 fps.
There are a total of eight transformer blocks that we may use to initialize our ControlNet. We chose the first six due to memory constraints. With more compute resources, one could potentially achieve higher quality control. We expose the number of transformer blocks to use in the ControlNet as a command line argument for convenience in our training code release.
@@ A.2 Additional details: mass understanding quantitative study
Each ball is subjected to forces of magnitude $0.125\cdot n$, where $n\in\{1,\dots,8\}$, with 10 videos generated per force value using different random seeds. To ensure experimental control, we utilize the GPT-Image-1 API to generate first frames and their variations, maintaining consistent initial ball positions and shapes across conditions. We automatically detect ball position using a Faster R-CNN model with a ResNet-50-FPN backbone (Ren et al., 2015), tracking the sports ball class (ID 37). We compute the distance traveled as the Euclidean distance in pixel space between detected bounding box centers in the first and last frames.
@@ A.3 Additional details: encoding strategy, point force
Appendix B Additional qualitative results
B.1 Failures and Limitations
@@
B.2 Extended comparison with physics simulation models
To demonstrate the point force model’s versatility, we curate a benchmark using first-frame images from prominent physics-in-the-loop papers: PhysDreamer (Zhang et al., 2024a), DreamPhysics Huang et al. (2024), MotionCraft (Aira et al., 2024), PhysGaussian (Xie et al., 2024), PhysGen (Liu et al., 2024b), PhysGen3D (Chen et al., 2025), Physics3D (Liu et al., 2024a), and PhysMotion (Tan et al., 2024). We apply our force prompting approach to these diverse scenarios, including poking plants (alocasia, ficus, bouquet of flowers), moving vehicles (boat in water, toy cars), and household objects (rocking horse). Our video results (see project webpage) illustrate that our purely neural method can handle the same visual scenarios almost as effectively as approaches requiring explicit physics simulation at inference time. These qualitative results are in line with our findings in Table 2.
Appendix C Additional emergent phenomena
C.1 Case study #1: Does the model enforce physical affordances?
@@
C.2 Case study #2: Does the model understand atomicity of objects?
@@
C.3 Case study #3: Does the model preserve cinematic effects of the original image?
@@
Appendix D Impact Statement
@@
Appendix E Survey Details and Instructions

### v2

@@ 6 Mass Understanding
Force-Mass Relationship Qualitative Study: We evaluate mass understanding across four benchmark tasks featuring geometrically identical objects with different implied masses. Our test scenarios includes ornaments (wooden versus cast iron), laundry baskets (empty versus filled with clothes), book stacks (one, two, or three books), and cube stacks (single versus double cube). To ensure experimental control, we utilize the GPT-Image-1 API to generate initial frames with variations where only the implied mass differs between conditions. Figure 6 presents some of these results, with demonstrating that lighter objects consistently travel farther when subjected to identical forces. This pattern remains robust across four random seeds. This behavior suggests an emergent understanding of mass-dependent physics in our force-prompted model. Other results are in the project webpage. Additionally, we find that the mass understanding behavior persists in the zero-shot multiple objects setting. We include these experimental details in Appendix B.2.
@@ Acknowledgments
We would like to thank Bill Freeman, Miki Rubinstein, Junyi Zhang, Junhwa Hur, Noah Fischer, Saining Xie, Calvin Luo, Shijie Wang, Koven Yu, Tian Yun, and Zilai Zeng for useful discussions. We would also like to thank the anonymous NeurIPS reviewers. This project was partially supported by Samsung. Our research was conducted using computational resources at the Center for Computation and Visualization at Brown University.
@@ A.1 Training hyperparameters
We train for 5000 steps with an effective batch size of 8 (using gradient accumulation over two steps) and saved model checkpoints every five hundred steps. We used bf16 mixed precision with tf32 support and initialized from the THUDM/CogVideoX-5b-I2V pretrained weights. ControlNet (zhang2023adding) was initialized from the first six transformer layers, for all three input channels, using a downscaling factor of eight, and a unit weight coefficient. We use the AdamW optimizer (adamw) (initial learning rate $1\times 10^{-5}$, $\beta_{1}=0.9$, $\beta_{2}=0.95$, maximum gradient norm 1.0) under a cosine-with-restarts learning-rate schedule (one cycle, 250 warm-up steps). We set our random seed to 42. Training videos contained up to 49 frames at 8 fps.
There are a total of 42 transformer blocks that we may use to initialize our ControlNet. We chose the first six due to memory constraints. With more compute resources, one could potentially achieve higher quality control. We expose the number of transformer blocks to use in the ControlNet as a command line argument for convenience in our training code release.
@@ A.2 Additional details: mass understanding quantitative study
Each ball is subjected to forces of magnitude $0.125\cdot n$, where $n\in\{1,\dots,8\}$, with 10 videos generated per force value using different random seeds. To ensure experimental control, we utilize the GPT-Image-1 API to generate first frames and their variations, maintaining consistent initial ball positions and shapes across conditions. We automatically detect ball position using a Faster R-CNN model with a ResNet-50-FPN backbone (ren2015faster), tracking the sports ball class (ID 37). We compute the distance traveled as the Euclidean distance in pixel space between detected bounding box centers in the first and last frames.
@@ A.3 Additional details: encoding strategy, point force
A.4 Additional details: comparison with PhysDreamer
We took six flower demo videos from the PhysDreamer teaser videos (Alocacia, Carnation, Orange Rose, Red Rose, White Rose, Tulip) and took a screenshot of a still frame from the video that did not have any force annotation on it. We passed these six frames, as well as the six equivalent force prompts (i.e. the same force vector from the original demo) into the Force Prompting model. We presented these videos side by side in the human study, which we served using Qualtrics. Both videos for a given scene were shown to the human annotator simultaneously, with left/right randomization.
Appendix B Additional details: zero-shot multiple forces
B.1 Multiple forces for multiple objects, benchmark
Our experiments confirm that the model successfully handles multiple simultaneous forces without requiring retraining. This capability emerges zero-shot by simply adding multiple Gaussian blobs to the control signal videos—one for each applied force. We include these videos in the project webpage. The 6 scenes we tested are:
  • An image with 2 Christmas tree ornaments
  • An image with 2 toy cars
  • An image with 2 vases, each with a flower in it
  • An image with 2 roses leaning diagonally, growing out of the ground
  • An image with 2 dandelions growing in a field
  • An image with 2 apples on separate branches of a tree
We generated these images using the GPT-Image-1 API. To test each image, we identified two directions where it would be reasonable to poke each object; for example, the dandelions can each be poked to the left and to the right. Then for each image, we used 4 different multiple-force force prompts, representing the $2\times 2$ different ways of combining the two different directions; for example, we can poke both dandelions to the left, poke both to the right, poke one to the right and the other to the left, then poke one to the left and the other to the right. We computed one video for each, using the same seed for all of them, with no cherry-picking. We found that $5/6$ of the videos in this multi-poke benchmark had perfect force adherence, and one of them had nearly perfect force adherence. The only failure case was in the two apples scene; when the left apple is poked to the right and the right apple is poked to the left, they move as if they were both poked to the right.
B.2 Multiple forces for multiple objects, mass understanding
We used the GPT-Image-1 API to construct images where two objects of different apparent masses are present. The first image contains an empty laundry basket on the left side of the frame, and a full laundry basket on the right side of the frame. The second image contains a book on the left side of the frame, and a stack of books on the right side of the frame. We passed both sets of images into the multi-poke Force Prompting inference pipeline and instructed the model to poke both sets of objects towards the middle of the frame with the same force magnitude. Across 8 different force magnitudes $\{0.125*i,i\in\{1,\dots,8\}\}$ we found that the lighter object moved much further, with the heavier item barely moving at all. We include videos on our project page.
Appendix C Additional qualitative results
C.1 Failures and Limitations
@@
C.2 Extended comparison with physics simulation models
To demonstrate the point force model’s versatility, we curate a benchmark using first-frame images from prominent physics-in-the-loop papers: PhysDreamer (zhang2024physdreamer), DreamPhysics huang2024dreamphysics, MotionCraft (aira2024motioncraft), PhysGaussian (xie2024physgaussian), PhysGen (liu2024physgen), PhysGen3D (chen2025physgen3d), Physics3D (liu2024physics3d), and PhysMotion (tan2024physmotion). We apply our force prompting approach to these diverse scenarios, including poking plants (alocasia, ficus, bouquet of flowers), moving vehicles (boat in water, toy cars), and household objects (rocking horse). Our video results (see project webpage) illustrate that our purely neural method can handle the same visual scenarios almost as effectively as approaches requiring explicit physics simulation at inference time. These qualitative results are in line with our findings in Table 2.
C.3 Training a unified model
We also train a unified model to learn both point force prompts and wind force prompts. We run inference on this joint model using our point force benchmark, as well as our wind force benchmark. Our findings:
  • More dynamic backgrounds: in many of the videos, the background moves more dynamically in a natural way. For example, for the apple tree, the surrounding leaves move more naturally after the apple is poked; in the video where a kid is sitting in the middle of a toy train track, the kid is moving more naturally while the train is moving around the track; and in the video with falling leaves in the forest with a woman sitting on a chair in the background, the woman in the chair moves more while the leaves are being blown.
  • Slightly less robust point force control: on some of the point force videos (e.g. the blueberry bush), the control signal is not respected.
We designed this experiment by sourcing 50% of the training data and control signals in each batch from the synthetic point force dataset, and the other 50% from the synthetic wind force dataset. We trained using the same architecture and number of training steps as the original model.
C.4 Scaling the dataset size
We trained from scratch a wind force model which uses half as many synthetic flag waving videos. The only variable that we changed was the dataset size; everything else (including the number of training steps and learning rate scheduler) remained the same. For this model, we found some additional failure cases. One failure case: for the image where a woman holds a sheet on the beach, the model hallucinates a bedsheet in the background, indicating that the model has likely memorized the “waving flag” pattern from the training dataset and is injecting it into the output video. A second failure case: the confetti’s response to the wind isn’t as convincing. For example, some of the confetti will blow in the wind’s direction but some will stay stationary. This indicates that the model hasn’t generalized properly, perhaps because of less training data.
Appendix D Additional emergent phenomena
D.1 Case study #1: Does the model enforce physical affordances?
@@
D.2 Case study #2: Does the model understand atomicity of objects?
@@
D.3 Case study #3: Does the model preserve cinematic effects of the original image?
@@
Appendix E Impact Statement
@@
Appendix F Survey Details and Instructions